OpenStudy (anonymous):

Jack bought 3 slices of cheese pizza and 4 sclices of mushroom pizza for $12.50. Grace bought 3 slices of cheese pizza and 2 slices of mushroom pizza for $8.50. What is the cost of 1 slice of mushroom pizza? A) $1.50 B) $2.00 C)$3.00 D)$3.50

OpenStudy (anonymous):

In this case, c=cheese pizza and m=mushroom pizza. (: For Jack: 3c+4m=12.50 For Grace: 3c+2m=8.50 Subtract the two equations above: 3c+4m=12.50 -(3c+2m=8.50) _________________ 2m=4 m=2 Answer is B) $2.00